OpenStudy (anonymous):

Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each?

OpenStudy (unklerhaukus):

\[T+B=56\]\[B>2\times T\]

OpenStudy (compassionate):

T + B = 56 B > 2 T, Now substitute, but you probably don't know what it means to substitute. So let's brake this down cause we're complete retards: 56-T = 2T+8 48 - T = 2T 48 = 3T T = 16 So it's actually really simple...

OpenStudy (compassionate):

T=16 t B= 56-16 = 40 And we're done...